OpenStudy (precal):

A right circular cylinder is such that the sum of the height and the circumference around the base is 30 cm. What is the radius that produces a maximum volume?

OpenStudy (precal):

@ganeshie8

OpenStudy (precal):

I believe this is optimization

OpenStudy (precal):

\[h+2 \pi r =30\]

OpenStudy (precal):

I am not sure if I can make this statement appreciate any help @zepdrix

OpenStudy (precal):

\[V=\pi r^2 h\]

OpenStudy (precal):

I wanted to use the first statement into the second statement, but not sure I am doing this correct. I know the final answer but I am trying to figure out the process

zepdrix (zepdrix):

Mmmm I think you're on the right track.\[\Large\rm V(r)=\pi r^2 (h)\]\[\Large\rm V(r)=\pi r^2 (30-2\pi r)\]\[\Large\rm V(r)=30\pi r^2-2\pi^2 r^3\]And then we want to differentiate with respect to...... r I guess?

zepdrix (zepdrix):

And then set our Volume derivative equal to zero. And look for r values that correspond to critical points of the volume function.

OpenStudy (precal):

ok here is a really dumb question on my part, when I take the derivative of the first term do I end up with 60 (pi)^2 r?

OpenStudy (precal):

pi is a constant and I am starting to confuse myself

zepdrix (zepdrix):

hehe

OpenStudy (precal):

or is it 60 pi r?

OpenStudy (precal):

glad you find this funny.........I have been going cross eye trying to do this problem

zepdrix (zepdrix):

(30pi) r^2 turns into 2(30pi) r or 60pi r yah the second one.

zepdrix (zepdrix):

sorry sorry >.< simmer down!

OpenStudy (precal):

no I am fine now........ok lets see if I can finish this problem

OpenStudy (precal):

how about that second term? with the pi^2?

OpenStudy (precal):

got it!!!! Finally ok so I got

OpenStudy (precal):

\[V ' = 60 \pi r - 6 \pi^2 r^2\]

OpenStudy (precal):

\[60 \pi r - 6 \pi^2 r^2=0\] \[6 \pi r = 0 \therefore r = 0\] \[10-\pi r = 0 \therefore r =\frac{ 10 }{ \pi }\]

zepdrix (zepdrix):

oo nice c:

OpenStudy (precal):

since the want the max it is 10/pi

OpenStudy (precal):

Thank You :)

OpenStudy (precal):

I really appreciate your help

zepdrix (zepdrix):

There was the other value, yes? r=0. But that minimizes I guess.

OpenStudy (precal):

yes but it wants the max

zepdrix (zepdrix):

cool c: